OpenStudy (anonymous):

Solve the equation. x - 9 = -6x + 5 Solve the equation. x - 9 = -6x + 5

OpenStudy (cicilybailey14):

x=2

OpenStudy (cicilybailey14):

Let's solve your equation step-by-step. x−9=−6x+5 Step 1: Add 6x to both sides. x−9+6x=−6x+5+6x 7x−9=5 Step 2: Add 9 to both sides. 7x−9+9=5+9 7x=14 Step 3: Divide both sides by 7. 7x 7 = 14 7

OpenStudy (anonymous):

yep

OpenStudy (yolo_king):

^^^^^^CORRECT^^^^^^^^^